|z-2|=2|z-1| then show that |z|^2= 4/3 real part of z using the properties of modulus and conjugate

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Vishesh
On squaring both the sides
Z² - 4Z + 4 = 4 ( Z² +1 -2Z)
Z² - 4Z + 4 = 4Z² + 4 -8Z
3Z² -4Z = 0
Z² = 4Z/3
Z² is a real number .
Hence it is 4/3 of real part of Z.
Regards
Arun (askIITians forum expert)